# Rethinking Epistemic and Aleatoric Uncertainty for Active Open-Set Annotation: An Energy-Based Approach

### v1

@@ 4.2 Performance Comparison
Figure 5: Ablation results on CIFAR-100 with a mismatch ratio of 40%. 1) The left validates the effectiveness of each method component. “Part 1” represents energy-based epistemic uncertainty and “Part 2” indicates energy-based aleatoric uncertainty. “mQP” denotes mean query precision across all AL rounds. 2) The middle assesses the sensitivity of the energy loss weight $\lambda_{e}$. “mDA” denotes mean detector test accuracy across all AL rounds. 3) The right shows the runtime comparison. The numbers on the bar chart correspond to mQP scores.
Figure 6: Ablation results for target precision $tP$ on CIFAR-10 (Left) and CIFAR-100 (Right). “MR” denotes mismatch ratio. “Best” indicates the top-performing method in the comparisons.
@@ 4.3 Ablation Studies
Effect of each component. Figure 5 (left) shows the ablation results to validate the effectiveness of each component in our method. Our proposed energy-based epistemic uncertainty (“Part 1”) utilizes information from labeled unknown class examples, leading to significant improvements in both recognition and test performance compared to free energy alone. The proposed energy-based aleatoric uncertainty (“Part 2”) performs poorly on its own, as aleatoric uncertainty is only meaningful in closed-set scenarios. However, when combined with “Part 1”, it shows significant improvement, validating the superiority of the entire framework. Additionally, removing the data-driven epistemic uncertainty score or the margin-based energy loss leads to a decline in performance, confirming their necessity.
@@ Acknowledgements
This work was supported by the National Key R&D Program of China (2020AAA0107000), NSFC (62222605), and the Natural Science Foundation of Jiangsu Province of China (BK20222012, BK20211517).
@@ Appendix A Uncertainty Quantification
    • In Figure 7, the “Bulldog” and “Arctic Wolf” exhibit significant feature differences from the “Dog” and “Wolf” in the training set, leading to higher epistemic uncertainty. After these examples are incorporated into model training, predictive performance on them improves, thereby reducing their epistemic uncertainty.
@@
    • In Figure 7, the “Seppala Siberian Sleddog” resembles the “Wolf” in appearance but belongs to the “Dog” class, leading to higher aleatoric uncertainty. Due to feature confusion, incorporating these examples into model training may not substantially improve performance or reduce their aleatoric uncertainty.
@@ Appendix E Additional Ablation Studies
Figure 8: Ablation results for $K$ in reverse k-NN on CIFAR-10 (Left) and CIFAR-100 (Right). “MR” denotes mismatch ratio. “Best” indicates the top-performing method in the comparisons.

### v2

@@ 4.2 Performance Comparison
Figure 5: Ablation results on CIFAR-100 with a mismatch ratio of 40%. 1) The left validates the effectiveness of each method component. “Part 1" represents energy-based epistemic uncertainty and “Part 2" indicates energy-based aleatoric uncertainty. “mQP" denotes mean query precision across all AL rounds. 2) The middle assesses the sensitivity of the energy loss weight $\lambda_{e}$. “mDA" denotes mean detector test accuracy across all AL rounds. 3) The right shows the runtime comparison. The numbers on the bar chart correspond to mQP scores.
Figure 6: Ablation results for target precision $tP$ on CIFAR-10 (Left) and CIFAR-100 (Right). “MR" denotes mismatch ratio. “Best" indicates the top-performing method in the comparisons.
@@ 4.3 Ablation Studies
Effect of each component. Figure 5 (left) shows the ablation results to validate the effectiveness of each component in our method. Our proposed energy-based epistemic uncertainty (“Part 1") utilizes information from labeled unknown class examples, leading to significant improvements in both recognition and test performance compared to free energy alone. The proposed energy-based aleatoric uncertainty (“Part 2") performs poorly on its own, as aleatoric uncertainty is only meaningful in closed-set scenarios. However, when combined with “Part 1", it shows significant improvement, validating the superiority of the entire framework. Additionally, removing the data-driven epistemic uncertainty score or the margin-based energy loss leads to a decline in performance, confirming their necessity.
@@ Acknowledgements
This work was supported by the Natural Science Foundation of Jiangsu Province of China (BK20222012) and the NSFC(U2441285, 62222605).
@@ Appendix A Uncertainty Quantification
    • In Figure 7, the “Bulldog" and “Arctic Wolf" exhibit significant feature differences from the “Dog" and “Wolf" in the training set, leading to higher epistemic uncertainty. After these examples are incorporated into model training, predictive performance on them improves, thereby reducing their epistemic uncertainty.
@@
    • In Figure 7, the “Seppala Siberian Sleddog" resembles the “Wolf" in appearance but belongs to the “Dog" class, leading to higher aleatoric uncertainty. Due to feature confusion, incorporating these examples into model training may not substantially improve performance or reduce their aleatoric uncertainty.
@@ Appendix E Additional Ablation Studies
Figure 8: Ablation results for $K$ in reverse k-NN on CIFAR-10 (Left) and CIFAR-100 (Right). “MR" denotes mismatch ratio. “Best" indicates the top-performing method in the comparisons.